y=2x
find four solutions for equations, write the solutions as ordered pairs.

Answer by jim_thompson5910(35256)

x=0, y=2x=2(0)=0
When x=0, y=0, giving us (0,0)


x=1, y=2x=2(1)=2
When x=1, y=2, giving us (1,2)


x=2, y=2x=2(2)=4
When x=2, y=4, giving us (2,4)


x=3, y=2x=2(3)=6
When x=3, y=6, giving us (3,6)